How many combinations can you make with 15 apples and 27 banana's?
Answer by Alan3354(69443)

How many combinations can you make with 15 apples and 27 banana's [sic]?
----------
Why do you have banana's, and not apple's?
-----------
If the apples and bananas (not banana's) are interchangeable, then it's
=1*2 = 2 possiblities
Either apple, banana or
banana, apple
-------
If each specimen is different, then it's
15*27 = 405